OpenStudy (anonymous):

please explain not just give answer What is the exact value of the expression 2the square root of 7. + the square root 14. + 4the square root of 7.? 4 the square root of 7. + the square root of 14. 6 the square root of 7. + the square root 14. 7 the square root of 28. the square root of 28.

jimthompson5910 (jim_thompson5910):

let x be the square root of 7 let y be the square root of 14 \[\Large 2\sqrt{7} + \sqrt{14} + 4\sqrt{7}\] \[\Large 2x + y + 4x\]

jimthompson5910 (jim_thompson5910):

simplify 2x+y+4x and then substitute the root values back in

OpenStudy (danjs):

For the common roots , addition can look like this \[a \sqrt{b} + c \sqrt{b} = (a+c)\sqrt{b}\]

OpenStudy (danjs):

For unlike roots, you can not add \[\sqrt{14}+\sqrt{7} = \sqrt{14}+\sqrt{7}\]

OpenStudy (anonymous):

im lost...

OpenStudy (danjs):

If you have the same number under the root sign, you can add them together. for example. \[4\sqrt{5}+\sqrt{3 }+6\sqrt{5} = (4+6)\sqrt{5} + \sqrt{3} = 10\sqrt{5}+\sqrt{3}\]

jimthompson5910 (jim_thompson5910):

Do you see how I replaced all the copies of "square root of 7" with x ?